OpenStudy (anonymous):

Calculate the following sum (100 values): + 11 13 15 ... 209

OpenStudy (ash2326):

@Social We have the series, let the sum be S S=11+13+15..............+207+209 I can also write sum as S=209+207.....+15+13+11 Do you follow this?

OpenStudy (anonymous):

i know it

OpenStudy (anonymous):

then? i do not remember

OpenStudy (ash2326):

okay, then let's add the two series, we'll add term by term 11+209=220 13+207=220 15+205=220 so we get 2S=220+220+220.....+220+220 Do you follow this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

can you find the sum from here? it's just a step

OpenStudy (anonymous):

11000

OpenStudy (ash2326):

correct :) did you understand the solution?

OpenStudy (anonymous):

yes... this is very clever it means that i will have the same result with different numbers

OpenStudy (anonymous):

many many years since school... i forgot them.. thank you!!

OpenStudy (ash2326):

Welcome :D